# Datacube segmentation via Deep Spectral Clustering

### v1

@@ 1 Introduction
The advent of Machine Learning (ML), and more specifically Deep Learning (DL), has revolutionised numerous fields, from computer vision to natural language processing [1]. In recent years, these powerful computational tools have also begun to reshape the field of applied physics.
@@
Deep Learning, a subset of ML, employs artificial neural networks with multiple layers (hence “deep”) to model and understand complex patterns. In the realm of applied physics, DL can be used to recognise patterns in large datasets, predict system behaviour, or control physical systems.
@@
One of the key advantages of Deep Learning in applied physics is its ability to process and learn from vast amounts of data, often surpassing traditional computational physics methods in both speed and accuracy. This is particularly useful in fields where experimental or simulation data is abundant, such as astrophysics and nuclear physics applied to cultural heritage.
@@
For instance, in astrophysics, the growing influx of data from advanced astronomical facilities calls for the use of DL algorithms, which excel in modeling complex patterns and handling large-scale data, often outperforming traditional data analysis methods (see [2, 3], and references therein). In the research field of galaxy formation and evolution, DL methods, such as convolutional neural networks (CNNs), have shown remarkable proficiency. For instance, recent studies [4, 5, 6] have employed these methods for more accurate and efficient galaxy morphological classification, a critical aspect in understanding galaxy evolution. Similarly, deep learning has revolutionized the study of gravitational lensing. For instance, the work [7] illustrates how deep neural networks can analyze Hubble Space Telescope images to measure galaxy mass distribution more precisely, thus enhancing our understanding of cosmic structures. The field of exoplanet discovery has also greatly benefited from DL (see e.g., [8, 9]). For instance, [10] have successfully applied neural networks to Kepler mission data, significantly improving the detection rate of new exoplanets. In cosmology, the analysis of the Cosmic Microwave Background has been revolutionized by DL (see, e.g., [11, 12]. and references therein), as highlighted in recent works [13, 14], which have extracted subtle cosmological signals from complex datasets more effectively than traditional methods. Moreover, the application of DL in transient astrophysical phenomena, such as fast radio bursts and supernovae, has enabled rapid data processing and analysis, crucial for timely scientific observations (e.g., [15, 16]). In summary, the growing field of deep learning in astrophysics is not just a technical advancement but a necessary evolution to keep pace with the ever-growing complexity and volume of astronomical data.
@@
In the field of nuclear physics applied to cultural heritage, DL can be used to analyze data from non-destructive testing techniques, such as neutron imaging or gamma-ray spectroscopy. For example, it can help identify the composition of ancient artefacts or detect hidden structures in historical buildings, providing valuable insights without damaging these precious objects. This involves the use of advanced algorithms to statistically analyse the data obtained from imaging using nuclear techniques. These algorithms can identify patterns and correlations in the data, providing valuable insights into the artwork’s composition, age, and condition. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30] (for other Machine learning approaches in Cultural Heritage, see [31], and references therein).
@@
Out of many deep learning methods, Deep Clustering is one approach that may be relevant for the topic at hand (for nice reviews of Deep Clustering, see [32, 33, 34], and references therein). Deep clustering is a branch of unsupervised learning that combines traditional clustering methods with deep learning.
@@
There are several approaches to deep clustering, including autoencoder-based methods, deep embedded clustering, and methods based on self-supervised learning. Autoencoder-based methods use an autoencoder to learn a compressed representation of the input data, and then perform clustering on the compressed data. Deep embedded clustering extends this idea by jointly optimizing the autoencoder and the clustering objective.
@@ 2 Methods: deep clustering with AutoEncoders
Having a well-trained architecture of that sort, starting from a datacube $\mathcal{I}_{x,y;e}$ of shape $(0pt,0pt,0pt)$, where the first two indices represent the $x-$ and $y-$pixel position, while $e$ is the spectral index, we can extract the $0pt\cdot 0pt$ $0pt-$dimensional spectra $X^{(i=1,\ldots 0pt\cdot 0pt)}$, pass them to the Encoder to get a lower-dimensional statistical representation of the data $\mu^{(i)}=\mathrm{Enc}[X^{(i)}]$, and perform a clustering onto those, to get a set of clusters $C_{I=1,\ldots N}$, and their barycenter $c_{I}=\frac{1}{\#C_{I}}\sum_{j|\mu^{(j)}\in C_{I}}\mu^{(j)}$. By using the Decoder is now possible to see how the barycenter maps in the spectral space, $\bar{X}_{I}=\mathrm{Dec}[c_{I}]$; furthermore, it is possible to create binary maps of clusters onto the image, by retaining the pixel-to-index relation, $(x,y)\leftrightarrow i$. We will show that those binary maps represents effectively unsupervised segmentation of the datacubes.
@@ 2.1 AutoEncoders Architecture
The first ingredient is an (AE) architecture [1, 37, 38, 39], based on either on plain Multi-Layer Perceptrons (MLP) with Drop-out layers, or Self-Normalising Neural Networks (SNN) [40]. The AutoEncoder class can be instantiated with a tunable number of layers, while the nodes are algorithmically fixed, starting from the Input size and dividing either by $2\cdot k$ (where $k$ is the layer index), or by $2^{k}$, up to the latent space dimension $n$. The Encoder and Decoder are instantiated with a specular architecture, i.e. the innermost layer has the same size as the outermost one, and so on inwards. The layers’ weights are initialized using the Kaiming-He initialization [41]. For the MLP module, the layer’s activation is a ReLU and the dropout probability can be tuned, while of course, the activation for the SNN module is a SELU. For a visual representation of an autoencoder, see figure 1.
@@ 2.2 Iterative K-Means
The iterative selection works as explained in Algorithm $1$For an overview of applications of unsupervised metric in K-Means clustering, see [46] and references therein..
@@
${}\mathrm{silhouette}[C_{I=1,\ldots k}]=\frac{1}{k}\sum_{I=1}^{k}\frac{1}{\#C_{I}}\sum_{i=1}^{\#C_{I}}\frac{b_{I}(i)-a_{I}(i)}{\max_{i}[a_{I}(i),b_{I}(i)]}\,,{}$
${}a_{I}(i)=\frac{1}{\#C_{I}-1}\sum_{j\in C_{I},i\neq j}\mathrm{dist}(i,j)\,,\quad b_{I}(i)=\min_{J\neq I}\frac{1}{\#C_{J}}\sum_{j\in C_{I}}\mathrm{dist}(i,j)\,,{}$
@@ 2.3 Deep Clustering
Assembling the ingredients introduced in the sections above, the auto-encoder architecture and the iterative K-Means clustering, it is possible to build a deep clustering auto-encoder, by defining the total loss as
@@
Notice that we need to have an implementation of the IKMeans clustering algorithm which does not break the back-propagation algorithm. In order to do so, we have written a PyTorch module for that, by ensuring adherence to the differentiable programming paradigm.
@@ 2.4 A Deep Clustering Variational AutoEncoder? Variational Deep Embedding
Additionally, we allow for either $\beta$ and $\gamma$ to be epoch-dependent, in order to tune dynamically during training how the MMD and the Silhouette terms affect the reconstruction, i.e. the loss we use is
@@ 3 Results: deep clustering synthetic datacubes
Since the work is an exploratory phase, we applied the developed architecture on two synthetic datasets (ordered alphabetically). The first is a synthetic dataset of astrophysical interest; the second, is a synthetic dataset of imaging with nuclear techniques on pictorial artworks.
@@ 3.1 Deep Clustering Astrophysical synthetic datacubes
To create a synthetic dataset of datacubes with astrophysical meaning, we started from synthetic 1D signals belonging to three classes: HII regions, planetary nebulae and shock regions. These classes represent three types of ionized gas nebulae found in galaxies, characterized by distinctive morphological features and ionization mechanisms. As a consequence, when observed, they produce optical spectra, which differ mainly in the relative intensities of their spectral lines. Therefore, traditional classification diagnostics are generally based on the differences in specific line ratios but do not exploit the information contained in their entire spectra, e.g., the Baldwin–Phillips–Terlevich (BPT) diagrams [66]. These techniques generally suffer from theoretical limits (see e.g. [67]), and often require human aid or a specific and complex analysis, thus they are less suitable for handling large datasets, obtained from current instruments with integral field units. This is why Machine Learning algorithms, capable of encoding the information from a wide sample of emission-line regions in the full spectra, are a valid choice.
We generate a collection of synthetic nebular optical spectra, based on photoionization models run with CLOUDY [68], withdrawn from the Mexican Million Models database (3MdB) [69]. We augment the number of models by combining them linearly using random coefficients. To create mock spectra starting from these models, we first select relevant emission lines to simulate, based on real data. In this case, we create spectra based on observations of the local galaxy M33 obtained with the Multi Unit Spectroscopic Explorer (MUSE) [70]. Hence, the selected emission lines are defined in the MUSE wavelength domain. Simulated emission lines are modeled with Gaussian profiles, whose fluxes are given by the model line intensities, and whose widths consist of two components: an instrumental width, caused by the finite spectral resolution of the telescope, and a physical one, given by the velocity dispersion in the nebulae, where the latter is chosen randomly from a range of possible values. Then, Gaussian noise (with standard deviations based on real data) is added to the spectra. Thus, for each model resulting from the linear combination of two of the ”original” database models, we can create multiple spectra with different noise and line widths.
@@ 3.1.1 Results:
We applied the trained model on the synthetic X-Ray Fluorescence (XRF) datacube obtained using ganX [27] on the RGB data of Figure 3 with threshold $0.2$, giving rise to 4 clusters.
Figure 4: Clustered datacube using the trained model. On the left, is the binary map of the cluster pixels. In the middle, the Reconstructed Signal averaged over the cluster; on the right, the “true” signal averaged over the cluster. Notice that the term ’XRF’ appearing here stands for ’signal’.
@@
Another relevant plot is Figure 5, which shows the energy-integrated data-cube, i.e. the gray-scale image (in viridis color scale) obtained integrating over the whole energy channel. The left image in Figure 5 represents the original, “true” [^2] Here, we refer to the “true” signal as the input signal of the neural network, while with ”reconstructed” we refer to the output of the network. spectral energy-integrated datacube $I_{i,j}=\sum_{e}I_{i,j;e}$. The middle image is the decoded energy-integrated datacube $D_{i,j}=\sum_{e}\mathrm{Dec}[\mathrm{Enc}[I_{i,j;e}]]$. It immediately shows noise reduction (or signal enhancement). The right image in Figure 5 represent the embedded energy-integrated datacube $E_{i,j}=\sum_{n}\mathrm{Enc}[I_{i,j;n}]]$, which further refine the readability.

### v2

@@ 1 Introduction
The advent of Machine Learning (ML), and more specifically Deep Learning, has revolutionised numerous fields, from computer vision to natural language processing [1]. In recent years, these powerful computational tools have also begun to reshape the field of applied physics.
@@
Deep Learning, a subset of ML, employs artificial neural networks with multiple layers (hence “deep”) to model and understand complex patterns. In the realm of applied physics, Deep Learning can be used to recognise patterns in large datasets, predict system behaviour, or control physical systems.
@@
One of the key advantages of Deep Learning in applied physics is its ability to process and learn from vast amounts of data, often surpassing traditional computational physics methods in both speed and accuracy. This is particularly useful in fields where experimental or simulation data is abundant, such as astrophysics or nuclear physics applied to Cultural Heritage.
@@
In astrophysics, the growing influx of data from advanced astronomical facilities calls for the use of Deep Learning algorithms, which excel in modeling complex patterns and handling large-scale data, often outperforming traditional data analysis methods (see [2, 3], and references therein). In the research field of galaxy formation and evolution, Deep Learning methods, such as Convolutional Neural Networks (CNNs), have shown remarkable proficiency. Recent studies [4, 5, 6] have employed these methods for more accurate and efficient galaxy morphological classification, a critical aspect in understanding galaxy evolution. Similarly, Deep Learning has revolutionized the study of gravitational lensing. For instance, the work [7] illustrates how deep neural networks can analyze Hubble Space Telescope images to measure galaxy mass distribution more precisely, thus enhancing our understanding of cosmic structures. The field of exoplanet discovery has also greatly benefited from Deep Learning (see e.g., [8, 9]). As an example, [10] have successfully applied neural networks to Kepler mission data, significantly improving the detection rate of new exoplanets. In cosmology, the analysis of the Cosmic Microwave Background has been revolutionized by Deep Learning (see, e.g., [11, 12]. and references therein), as highlighted in recent works [13, 14], which have extracted subtle cosmological signals from complex datasets more effectively than traditional methods. Moreover, the application of Deep Learning in transient astrophysical phenomena, such as fast radio bursts and supernovae, has enabled rapid data processing and analysis, crucial for timely scientific observations (e.g., [15, 16]). In summary, the growing field of Deep Learning in astrophysics is not just a technical advancement but a necessary evolution to keep pace with the ever-growing complexity and volume of astronomical data.
@@
In the field of nuclear physics applied to Cultural Heritage (for a nice introduction to the subject, see, e.g. [17, 18, 19, 20, 21, 22, 23, 24] and references therein), Deep Learning can be used to analyze data from non-destructive testing techniques, such as neutron imaging or gamma-ray spectroscopy. For example, it can help identify the composition of ancient artifacts or detect hidden structures in historical buildings, providing valuable insights without damaging these precious objects. This involves the use of advanced algorithms to statistically analyse the data obtained from imaging using nuclear techniques. These algorithms can identify patterns and correlations in the data, providing valuable insights into the artwork’s composition, age, and condition. [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38] (for other Machine Learning approaches in Cultural Heritage, see [39], and references therein).
@@
Out of many Deep Learning methods, Deep Clustering is one approach that may be relevant for the topic at hand (for nice reviews of Deep Clustering, see [40, 41, 42, 43, 44, 45], and references therein). Deep Clustering is a branch of unsupervised learning that combines traditional clustering methods with Deep Learning.
@@
There are several approaches to Deep Clustering, including autoencoder-based methods [47, 48], deep embedded clustering [49, 50], and methods based on self-supervised learning [51, 52, 53, 54]. Autoencoder-based methods use an autoencoder [55] to learn a compressed representation of the input data, and then perform clustering on the compressed data. Deep embedded clustering extends this idea by jointly optimizing the autoencoder and the clustering objective.
@@ 2 Methods: Deep Clustering with AutoEncoders
Having a well-trained architecture of that sort, starting from a datacube $\mathcal{I}_{x,y;e}$ of shape $(0pt,0pt,0pt)$, where the first two indices represent the $x-$ and $y-$pixel position, while $e$ is the spectral index, we can extract the $0pt\cdot 0pt$ $0pt-$dimensional spectra $X^{(i=1,\ldots 0pt\cdot 0pt)}$, pass them to the Encoder to get a lower-dimensional statistical representation of the data $\mu^{(i)}=\mathrm{Enc}[X^{(i)}]$, and perform a clustering onto those, to get a set of clusters $C_{I=1,\ldots N}$, and their barycenter $c_{I}=\frac{1}{\#C_{I}}\sum_{j|\mu^{(j)}\in C_{I}}\mu^{(j)}$. By using the Decoder it is now possible to see how the barycenter maps in the spectral space, $\bar{X}_{I}=\mathrm{Dec}[c_{I}]$; furthermore, it is possible to create binary maps of clusters onto the image, by retaining the pixel-to-index relation, $(x,y)\leftrightarrow i$. We will show that those binary maps effectively represent unsupervised segmentation of the datacubes.
@@ 2.1 AutoEncoders Architecture
The first ingredient is an AutoEncoder (AE) architecture [1, 56, 57, 58], based on either on plain Multi-Layer Perceptrons (MLP) with Drop-out layers, or Self-Normalising Neural Networks (SNN) [59]. The AutoEncoder class can be instantiated with a tunable number of layers, while the nodes are algorithmically fixed, starting from the Input size and dividing either by $2\cdot k$ (where $k$ is the layer index), or by $2^{k}$, up to the latent space dimension $n$. The Encoder and Decoder are instantiated with a specular architecture, i.e. the first Encoder layer has the same size as the last Decoder one, and so on inwards. The layers’ weights are initialised using the Kaiming-He initialisation [60]. For the MLP module, the layer’s activation is a ReLU and the dropout probability can be tuned, while of course, the activation for the SNN module is a SELU. For a visual representation of an autoencoder, see figure 1.
@@ 2.2 Iterative K-Means
The iterative selection works as explained in Algorithm $1$For an overview of applications of unsupervised metric in K-Means clustering, see [65] and references therein. (dubbed IKMeans).
@@
${}\mathrm{silhouette}[C_{I=1,\ldots k}]=\frac{1}{k}\sum_{I=1}^{k}\frac{1}{\#C_{I}}\sum_{i=1}^{\#C_{I}}\frac{b_{I}(i)-a_{I}(i)}{\max_{i}[a_{I}(i),b_{I}(i)]}\,,{}$ ${}a_{I}(i)=\frac{1}{\#C_{I}-1}\sum_{j\in C_{I},i\neq j}\mathrm{dist}(i,j)\,,\quad b_{I}(i)=\min_{J\neq I}\frac{1}{\#C_{J}}\sum_{j\in C_{I}}\mathrm{dist}(i,j)\,,{}$ (2)
@@ 2.3 Deep Clustering
Assembling the ingredients introduced in the sections above, the autoencoder architecture and the iterative K-Means clustering, it is possible to build a Deep Clustering autoencoder, by defining the total loss as
@@
It is important to notice that the implementation of the IKMeans clustering algorithm has to be written in a way that it does not break the back-propagation algorithm. We implemented IKMeans as a PyTorch module, which ensures adherence to the differentiable programming paradigm.
@@ 2.4 A Deep Clustering Variational AutoEncoder? Variational Deep Embedding
Additionally, we allow for either $\beta$ and/or $\gamma$ to be epoch-dependent, in order to dynamically tune how the MMD and the Silhouette terms affect the reconstruction during the training. The loss we use is
@@
2.5 Training vs Evaluation settings
Before moving on into the topic of synthetic datasets generation, and how the Deep Embedding models are trained and thus evaluated, we discuss in depth the logical steps necessary to get from a Deep (Variational) Embedding of 1D spectra to a 3D datacube unsupervised segmentation algorithm.
The model is trained on 1D spectra, i.e. the pixel-wise content of a standard datacube. In fact, we refer to datacube as the data structure $I_{i,j;e}$ described by three indices, $\{i,j;e\}$; $(i,j)$ indices refer to the pixel position along $X$ and $Y$ image axes, while $e$ is the energy channel. In analogy, an RGB image is a datacube with three possible energy channels, corresponding to the Red, Blue and Green pixel content. So, the RGB equivalent of the training process is to train a 1D model to embed single RGB triples into a different, low dimensional metric space; here, the spectra are multi-channels [^2] They should actually be continuous, representing actual spectra (in the visible/near-infrared region for the Astrophysical case, in the X-ray region for the Cultural Heritage case); their discrete nature is of course given by the Analog-to-Digital converter which bins the data into discretised digital values., living in a very high dimensional space ($\mathbb{N}^{512}$ or $\mathbb{N}^{1024}$ once rebinned).
Notice that the model does not learn from images (i.e., it does not have access to geographical information, like pixels proximity or location); it is trained over (synthetic) 1D signals.
The training of the Deep Neural Network model is done using mini-batch gradient descent (with Adam [84] optimiser). The batch size is kept quite large ($512$ $1D$ spectra per batch), in order to have a meaningful per-batch clustering; for each cluster, an iterative KMeans clustering is performed, to find a per-batch optimum number of clusters, in an overall fixed range of possible cluster numbers (which are training hyperparameters).
The fact that IKMeans is a distance-based clustering algorithm should force the Encoder model to map spectra into a well-behaved low dimensional space, which has to be metric (where the distance map inducing the metric is the same used in the IKMeans algorithm).
Furthermore, for each hyperparameter set training trial, a ReduceLROnPlateau learning rate scheduler is used, in its standard PyTorch implementation. We notice that our training is unsupervised and is based solely on 1D signals (not images, at this level).
After having trained the best model for the task, we try to assess the quality of the trained network. To do so, we generate a synthetic datacube, starting from a seed RGB image (how this is done will be explained in more depth in the following sections); we then flatten the image (thus seeing it as a collection of spectra), pass it to the trained model, get the embedded version (by reverse the flattening) and related clustering. We will show in the following that, as an emergent phenomenon, we experience a sort of semantic segmentation of the datacube in this process.
2.5.1 Alternative approaches: Deep Embedding vs Dimensional Reduction:
The emergence of a semantic segmentation through clustering in lower dimensional embedded space can also be obtained using other Dimensional Reduction algorithms, such as Principal Component Analysis (PCA) [85, 86] (for a modern introduction to the subject, see [87] and references therein) or Manifold Learning algorithms, such as t-distributed stochastic neighbour embedding (t-SNE) [88, 89] or Uniform Manifold Approximation and Projection (UMAP) [90] (for a recent review on manifold learning and non-linear dimensional reduction, see [91, 92, 93, 94, 95] and references therein).
While applying dimensional reduction algorithms to a spectral datacube, followed by a clustering algorithm is a viable method to get an emergent datacube semantic segmentation, this approach is rather different from the one we propose here.
Firstly, a Deep Embedding model, having a Decoder network, is invertible. While PCA is itself invertible, being a linear mapping, Manifold Learning such as t-SNE and UMAP are inherently difficult to invert.
Furthermore, t-SNE and UMAP (as well as other non-linear dimensional reduction algorithms) have a computational cost that scales rapidly with vector dimensions; this means that, usually, we need to perform a preliminary PCA reduction to an intermediate dimensional space [89] [^3] In [90], instead of PCA, a preliminary Spectral embedding is performed by considering the 1-skeleton of the global fuzzy topological representation as a weighted graph. They show that, empirically, UMAP is faster than t-SNE while scaling with ambient dimension; still, the computational cost does increase hugely, and non-linearly, even though more slowly.. Thus, while Deep Embedding models need a slow train phase, which may be slower compared to the other methods, it is fast in its inference phase.
Additionally, the embedding space may not have a metric meaning for non-linear dimensional reduction; this latter point may be relevant, because it means that distances in the embedded, low dimensional space are not proxies for similarity (i.e., nearby points are not necessarily more similar, nor far away points are necessarily more different).
PCA, as well as other linear dimensional reduction methods, are incapable of capturing complex, non-linear relations among data points, due to their linear nature. This also implies that PCA-reduced datasets may be difficult to cluster properly.
Finally, the crucial difference is that, while the deep embedding model is trained on a dataset formed as a collection of 1D spectra, and then applied onto datacubes (seen as a collection of spectra) at inference, the other models do need to be trained for each image (to have a well tailored behaviour).
In this sense, the two approaches are rather different:
  1. 1.
    in Deep Clustering, we train on an ad-hoc dataset and apply it on datacubes, and see what happens.
  2. 2.
    In dimensional reduction + clustering, we fit-and-predict on per-image basis and see what happens.
The two methods are different and, in some cases, may be complementary, since they fulfil different goals.
Finally, if we train a Deep Variational Embedding model, we add a generative spin, which cannot, in any case, be obtained in the Dimensional Reduction case.
In Table 1 we summarise these differences.
Deep Embedding PCA Manifold Learning
Invertible
Pure high dimensionality origin
Fast inference
Fast training /
Metric embedding space
Captures non-linear relations
Clustering-apt
Image independent
Generative /
Table 1: Summary of differences between Deep Embedding, PCA and Manifold Learning algorithms.
@@ 3 Results: Deep Clustering synthetic datacubes
Since the work is an exploratory phase, we applied the developed architecture on two synthetic datasets (ordered alphabetically). The first is a synthetic dataset of astrophysical interest; the second, is a synthetic dataset of imaging obtained by nuclear techniques on pictorial artworks.
@@
All the datasets, trained models, and produced code is publicly available online. See Section 5 for more details.
@@ 3.1 Deep Clustering Astrophysical synthetic datacubes
To create a realistic synthetic dataset of datacubes with astrophysical meaning, we started from synthetic 1D signals belonging to three classes: HII regions, planetary nebulae and shock regions. These classes represent three types of ionized gas nebulae found in galaxies, characterized by distinctive morphological features and ionization mechanisms. As a consequence, when observed, they produce optical spectra, which differ mainly in the relative intensities of their spectral lines. Therefore, traditional classification diagnostics are generally based on the differences in specific line ratios but do not exploit the information contained in their entire spectra, e.g., the Baldwin–Phillips–Terlevich (BPT) diagrams [97]. These techniques generally suffer from theoretical limits (see e.g. [98]), and often require human aid or a specific and complex analysis, thus they are less suitable for handling large datasets, obtained from current instruments with integral field units. This is why Machine Learning algorithms, capable of encoding the information from a wide sample of emission-line regions in the full spectra, are a valid choice.
We generate a collection of synthetic nebular optical spectra, based on photoionization models run with CLOUDY [99], withdrawn from the Mexican Million Models database (3MdB) [100]. We augment the number of models by combining them linearly using random coefficients. To create mock spectra starting from these models, we first select relevant emission lines to simulate, based on real data. In this case, we create spectra based on observations of the local galaxy M33 obtained with the Multi Unit Spectroscopic Explorer (MUSE) [101]. Hence, the selected emission lines are defined in the MUSE wavelength domain. Simulated emission lines are modeled with Gaussian profiles, whose fluxes are given by the model line intensities, and whose widths consist of two components: an instrumental width, caused by the finite spectral resolution of the telescope, and a physical one, given by the velocity dispersion in the nebulae, where the latter is chosen randomly from a range of possible values. Then, Gaussian noise (with standard deviations based on real data) is added to the spectra. Thus, for each model resulting from the linear combination of two of the “original” database models, we can create multiple spectra with different noise levels and line widths. This process is based on solid, realistic physical assumptions and includes typical noise distributions and specific instrumental effects, ensuring the resulting spectra are highly realistic.
@@ 3.1.1 Results:
We applied the trained model on the synthetic datacube obtained using ganX [35] on the RGB data of Figure 3 with threshold $0.2$, giving rise to 4 clusters.
Figure 4: Clustered datacube using the trained model. On the left, is the binary map of the cluster pixels. In the middle, the Reconstructed Signal averaged over the cluster; on the right, the “true” signal averaged over the cluster.
@@
Another relevant plot is Figure 5, which shows the energy-integrated data-cube, i.e. the gray-scale image (in viridis color scale) obtained integrating over the whole energy channel. The left image in Figure 5 represents the original, “true” [^4] Here, we refer to the “true” signal as the input signal of the neural network, while with “reconstructed” we refer to the output of the network. spectral energy-integrated datacube $I_{i,j}=\sum_{e}I_{i,j;e}$. The middle image is the decoded energy-integrated datacube $D_{i,j}=\sum_{e}\mathrm{Dec}[\mathrm{Enc}[I_{i,j;e}]]$. It immediately shows noise reduction (or signal enhancement). The right image in Figure 5 represent the embedded energy-integrated datacube $E_{i,j}=\sum_{n}\mathrm{Enc}[I_{i,j;n}]$, which further refine the readability.
3.1.2 Check against RGB clustering:
Since, as mentioned in Section 2, the datacube segmentation is an emergent result from clustering in latent space, and comes from a completely unsupervised approach, there is no direct evaluation check with standard semantic segmentation loss and measures, such as the Tversky Index (see, e.g., [102, 103, 104], and references therein).
Nervetheless, since we are discussing the application of synthetic datacubes, obtained starting from an RGB image (Figure 3), we can perform the same Iterative K-Means clustering approach to the RGB image, and use it as a baseline to compare it to the clustering obtained with the model.
This is a check that cannot be done in a real scenario, since
  1. 1.
    we may not have an RGB image;
  2. 2.
    RGB clustering may not be able to capture the relevant information;
Nevertheless, since our synthetic datacube generation pipeline relies on information extracted from the RGB image via a clustering process, we may use it as a proxy for a benchmark.
Figure 6: Deep Clustering Segmentation vs RGB Clustering Deep Embedding model on Astrophysical synthetic datacube.
In Figure 6 we report the result of this check, done cluster-by-cluster. The columns are, from left to right: the black-and-white image of the cluster mask (i.e., pixels are white if they do belong to the cluster, and black if they don’t) as obtained from the Deep Embedding Model; the same, but obtained from the RGB clustering; the image of mislabelled pixels (i.e., pixels appears red if they are mislabelled and blue if they are correctly labelled); the confusion matrix. On the confusion matrix is reported the number of counts (from left to right, from top to bottom: true positives, false positives, false negatives, true negatives). In the confusion matrix image subtitle is reported the computed Tversky Index. The Tversky Index (TI) formula is
$\mathrm{TI}=\frac{\mathrm{tp}}{\mathrm{tp}+\alpha\,\mathrm{fn}+\beta\,\mathrm{fp}}\,,$ (7)
we use the parameter set $\alpha=0.7,\beta=0.3$ as in [105].
Furthermore, in the Figure subtitle we reported the macro averages for Accuracy, Precision, Recall, F1 score and Tversky Index.
Figure 7: Multi-class Confusion Matrix of Deep Clustering Segmentation vs RGB Clustering on the synthetic Astrophysical Datacube.
In Figure 7 we report the multi-class confusion matrix; on each element of the confusion matrix is reported the counts normalised to the whole population. In the Figure subtitle is reported the Weighted average of Accuracy, Recall, Precision, and F1 score.
Notice that the result obtained from the datacube has a large count in the square (true label, predicted label) $=(4,1)$; this, again, is due to the fact that the network has not seen any synthetic background signal during training, and yet it was still capable of cluster (most of) the background pixels.
In B we report a test to show how the trained model behaves under the degradation of the signal (i.e., by adding noise such that the Peak-signal-to-Noise is reduced).